OpenStudy (anonymous):

What is the center and radius of the circle with the given equation? x^2+y^2+8x-4y=-11

OpenStudy (anonymous):

x² - 8x (leave a gap) + y² + 4y (leave a gap) = -11 to fill in the gaps, you complete the square divide coefficient of x by 2 and then square it. First gap: (8/2)² = 16 Second gap (4/2)² = 4 You have to add these numbers to both sides. x² - 8x + 16 + y² + 4y + 4 = -11 + 16 + 4 Now factor the expressions on the left and add the numbers on the right. (x - 4)² + (y + 2)² = 9 The center is (4, -2) and the radius is SQRT(9) = 3.

OpenStudy (anonymous):

There you goo :)

OpenStudy (anonymous):

Can i get a medal

OpenStudy (anonymous):

thank you!